a body of mass 20kg is placed on rough horizontal surface having mew is equal to 0.2 .a force of100n is applied on the body. calculate distance traveled by body in 10 sec

Grade:11

2 Answers

deepak
6 years ago
answer is so simple the max. friction is umg =0.2x20x10=40 
so net force acting is 100-40=60
so net acceleration is 60/20=3
so intial velocity =0
s=1/2at2=150m
 
Vikas TU
6 years ago
Dear Student,
Prefer this step by step method:
frictional force=µN=0.2 x 200=40N (Because N = Normal reaction = weight of the body = 200N)

Net force=100-40 = 60N
acceleration = Net force/mass = 60/20 = 3m/s2

S=ut+at2/2 (u=0, t=10s, a=3m/s2)
=>S=3x100/2
=150m.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)